# Interplay between strain and size quantization in a class of topological insulators based on inverted-band semiconductors

### v1

@@ Abstract
We consider surface states in semiconductors with inverted-band structures, such as $\alpha$-Sn and HgTe. The main interest is the interplay of the effect of a strain of an arbitrary sign and that of the sample finite size. We clarify the origin of various transitions which happen at a given strain with the change of the sample thickness, in particular the transition between the Dirac semimetal and quasi-3D (quantized) topological insulator. We compare our results with the ones recently published in the literature. For the k-p Kane model we derive effective boundary conditions in the case when the direct band materials form high barriers for the carriers of the inner inverted-band semiconductor (for example, CdTe/HgTe/CdTe and CdTe/$\alpha$-Sn/CdTe cases). We show that in this case the BCs have an universal and simple form which does not depend on the order of different non-commuting operators in the Hamiltonian. Even in the limit of very high barriers the BCs do not reduce to the trivial zero form, but contain the information about the asymmetry of the offsets in the conduction and valence bands. These boundary conditions allow to investigate the realistic case of finite mass of the heavy hole band, and to compare the results obtained within the Kane and the Luttinger models.
@@ I Introduction
Describing the electronic band structure of gapless semiconductors has long been an interesting problem because of the rich physics at their interfaces. Gapless semiconductors, like HgTe (or $\alpha$-Sn), have inverted band structures where the s-like $\Gamma_{6}$ ($\Gamma_{7-}$) and p-like $\Gamma_{8}$ ($\Gamma_{8+}$ ) bands switch order in energy due primarily to strong spin-orbit coupling. Groves ; Bloom ; Gelmont Crystalline symmetry enforces a four-fold degeneracy at the $\Gamma_{8}$-point, effectively closing the bandgap. Because of the band inversion, interfaces with vacuum or with direct band-gap semiconductors having ordinary band ordering can produce new types of interface (surface) states, namely Dyakonov-Khaetskii (DK) states Dyak and Volkov-Pankratov (VP) states Volkov ; Cade ; Suris ; Volkov1 . Moreover, these surface states have nondegenerate spin-texture, making them interesting for spintronics and optoelectronics applications. Ando Though these surface states were postulated in the 1980s, their properties remain consistent with what can be expected from topological phases of materials. Hasan ; Qi ; Bernevig ; Kane ; Dai ; Khaetskii ; Pankratov Experimentally, these interface states can be revealed by angle- and spin-resolved photoemission spectroscopy Xia ; Hsieh and magneto-transport measurements. Brune ; Kozlov ; Checkelsky ; Analytis ; Ren ; Xia1 ; Pan
@@
The energy dispersion of interface states becomes more complex for many practical samples of a thin topological insulator deposited onto a topologically trivial semiconductor or insulator. Under these circumstances, both finite-size effects and epitaxial strain can significantly modify the electronic band structure, causing transitions between different topological states. Because future devices employing topological materials will likely be made using thin films, the interplay of the finite-size effects and strain within a topological insulator is an important and interesting problem. In the current literature, there is a broad range of opinions about the origin and exact conditions for topological transitions of thin-film TIs Ohtsubo ; Coster ; Anh . The main goal of this work is to understand the physics of the series of transitions that happen at a given strain value with changing sample width.
@@
We consider here a model system comprised of a gapless semiconductor with an inverted band structure (e.g. HgTe or $\alpha$-Sn) of finite-width $d$ sandwiched between layers of a direct-gap (DG) semiconductor (e.g. CdTe or InSb), see Fig. 1. In section II, we present the Hamiltonian of the Kane model which includes also the contributions from the far bands; thus the HH band acquires a finite mass, which allows one to describe the physics applicable for realistic situations. In section III we develop the necessary boundary conditions (BCs) for a proper solution under finite thickness. In the case when the direct-band materials form high barriers for the carriers of the inner inverted-band semiconductor the BCs have an universal and simple form, thereby offering simple analytical treatment and more physical insight than achieved previously. In section IV we use these BCs within the Kane model with strain and finite HH mass to describe the interface between the inverted-band and direct-band semiconductors focusing on the energies of the DK surface states Dyak ; Khaetskii . This problem is important because of the necessity to clarify existing ARPES data. In section V we introduce finite-size effects and show how they split the VP states and discuss how it relates to possible interpretations of experimental data (for example, ARPES). Finally, in section VI, we describe the interplay between strain and finite-size effects for cases of biaxially compressive and tensile strain.
@@
Figure 1: Left: Configuration of the structure under consideration. A gapless semiconductor (1) of thickness $d$ is bounded on parallel sides by a direct-gap semiconductor (2). A Cartesian coordinate system is placed such that the $yz$-plane is within one interface where $x=0$. Right: Energy band diagram of the structure. The parameter $\eta$ describes the asymmetry of the offsets in the conduction and valence bands.
@@ II Material Configuration and Hamiltonian
III Boundary conditions
III.1 High-energy barriers between inverted and direct materials
The wave functions (spinors) in all regions have the form
Here $u$ corresponds to the electron contribution, while $v$ and $w$ describe the light-hole and heavy-hole contributions, respectively. We will consider further (except subsection III, C) the case of high-energy offsets in the conduction and valence bands where the energy gap of the direct barrier material is large compared to the energy gap of the inverted-band material. Using this condition we find the following decaying solutions in the barrier region, $x>0$, for the heavy and light particles:
$\Psi_{h}=e^{-\tilde{\kappa}_{h}x}\left(\begin{array}[]{ll}0&\\ 0&\\ 1\end{array}\right),\,\,\,\Psi_{l}=e^{-\tilde{\kappa}_{l}x}\left(\begin{array}[]{cc}-i\sqrt{\frac{1-\eta}{\eta}}&\\ 1&\\ 0\end{array}\right)$ (20)
Function Eq.(13) in the barrier region is a linear superposition of the functions Eq.(20). Here $\tilde{\kappa}_{l}=\sqrt{3\eta(1-\eta)/2}E_{gb}/P$, and $\tilde{\kappa}_{h}=\sqrt{(1-\eta)E_{gb}/(\gamma-2\tilde{\gamma})_{b}}$. The parameter $\eta$ ($10<\eta<$ 1) describes the asymmetry of the offsets in the conduction and valence bands, see Fig. 1. The edges of the conduction and valence bands in the barrier region are $E_{cb}=\eta E_{gb}$ and $E_{vb}=(\eta-1)E_{gb}$, where $E_{gb}>0$ is the energy gap of the direct-gap (barrier) material. The zero of the energy coincides with the middle of the $\Gamma_{8}-\Gamma_{6}$ gap of the inverted material. The case $\eta=1/2$ corresponds to symmetric offsets.
III.2 Boundary conditions
While deriving the boundary conditions, we will consider the lowest-order effects in small $\gamma,\tilde{\gamma}\ll P^{2}/E_{g}$, where $E_{g}>0$ is the $\Gamma_{8}-\Gamma_{6}$ energy gap in the inverted-band material. Consequently, one takes into account the finite values of $\gamma$ and $\tilde{\gamma}$ only in the heavy holes spectrum, thus the flat HH band acquires a finite mass. Assuming the high barrier condition, one obtains the following effective boundary conditions for the wave function inside the gapless material 1, see Fig. 1:
$\left(u+iv\sqrt{\frac{1-\eta}{\eta}}\right)|_{x=0}=0;\,\,\,w|_{x=0}=0$ (21)
We note that at the opposite boundary ($x=-d$) the sign in front of the imaginary unity in Eq. (21) should be changed for the negative one. By the effective boundary conditions we mean the following. With high barriers condition we can consider the solution of the Schrodinger equation only in the inner region 1, all the information about the barrier regions enters the problem through the BCs presented by Eq. (21). In this way we avoid the usual troubles Bastard related to the choice of the order of different noncommuting operators in the Hamiltonian, Eq. (8). It is remarkable that even in the limit of very high barriers the BCs do not reduce to the trivial zero form ($u=0,v=0,w=0$), but contain the information about the asymmetry of the offsets in the conduction and valence bands. It is also important to mention that while considering the surface states problem with high-energy barriers and for energies near the $\Gamma_{8}$ point (i.e. much smaller than the energy gap $E_{g}$) one can neglect the contribution from the electron $\Gamma_{6}$ band. Then the boundary conditions within the Luttinger model reduce to $v=0,w=0$, i.e. to zero values of the light-hole and heavy-hole components Dyak ; Dyak1 .
III.3 Location of the Dirac point of the Volkov-Pankratov states (a single interface)
For arbitrary values of the gap energies, $E_{g}>0,E_{gb}>0$, in the inverted and direct (barrier) materials, and arbitrary offsets in the conduction and valence bands, the exact formula for the location of the Dirac point of the Volkov-Pankratov states Volkov reads
Eq. (22) is obtained from the boundary conditions (continuity of all components of spinors found at $k_{y}=0$ in the inverted-band and direct-gap materials, while neglecting the $\gamma,\tilde{\gamma}$ terms).
Some ARPES experiments Engel show that the DP lies in close proximity to the $\Gamma_{8}$ point. This location can be obtained from Eq.22 in the case when $E_{gb}\gg E_{g}$ and $\eta\to 1$. $\eta\to 1$ means that the offset in the conduction band is much larger than in the valence band. Then at $\Delta=0$ one obtains $\epsilon_{DP}=E_{g}/2$, i.e. at the energy of the $\Gamma_{8}$ point of the inverted material.
If, however, one considers the case of $\alpha$-Sn/InSb interface Engel , then using the values of offsets from Ref. Cardona, and assuming a zero strain, one obtains the location of the DP to be rather closer to the middle of the $\alpha$-Sn gap, which differs from the interpretations Engel of the ARPES measurements.
IV Kane model with strain and finite heavy-hole mass
@@
The surface state spectrum near the $\Gamma_{8}$ energy for the realistic case of the finite heavy-hole mass is interesting because of the necessity to clarify existing ARPES data Engel . It is especially important to exclude or confirm the possibility of forming the twinning of Dirac cones,Seixas ; Balatsky which we set out to do here for our particular system.
@@
Two types of surface branches are found within the Kane model. The parabolic branch (called DK1 state) Dyak ; Khaetskii starts from the HH band and ends at the LH band. The Volkov-Pankratov states lie within the $\Gamma_{8}-\Gamma_{6}$ gap Volkov . The DK1 branch in the presence of strain was studied in Ref. Khaetskii, for the idealized case of a flat heavy-hole band. Here we investigate this problem at $\Delta>0$ (biaxial, in-plane tensile) and $\eta=1/2$ for a small ratio $\beta$ of the light and heavy hole masses given by $\beta=3\gamma_{h}E_{g}/2P^{2}\ll 1$, $\gamma_{h}=(\gamma-2\tilde{\gamma})$ We will use the perturbation approach with respect to $\beta\ll 1$.
@@
Using the boundary conditions in Eq. (21), one obtains the folowing equation for the surface states spectra (for the symmetric case of $\eta=1/2$):
@@
The quantity $B$ in Eqs. (29,32) is defined as $B=\epsilon+\Delta/2-E_{g}/2$. The equation for the time-reversed branch of the states is obtained by replacing $k_{y}$ by $-k_{y}$ in Eq. (32).
@@
where $\epsilon_{HH}=E_{g}/2-\Delta/2$ is the energy of the bottom of the HH band, and $\epsilon_{DP}=\Delta/4$ is the location of the Dirac point of the Volkov-Pankratov states (at $k_{y}=0$). We should mention that in contrast to the $\Delta=0$ case, for a finite positive strain there are two DK1 branches corresponding to the opposite spins for a given value of $k_{y}$. The one corresponding to the spin-up state (for $k_{y}>0$) exists only up to the critical value of $k_{y}=k_{\Delta}=\sqrt{6}\Delta/P$ (assuming that $\Delta\ll E_{g}$), where it merges with the heavy-hole band, see Fig. 2 and the discussion below.
For completeness we also present here (at $\beta=0$ ) the exact spectra of two V-P surface branches that lie within the $\Gamma_{8}-\Gamma_{6}$ energy gap (and which we derived using the formulas from Ref. Khaetskii, ):
@@
Note that Eq. (34) is the generalization for the case of $\Delta\neq 0$ of the equations of Ref. Volkov1, . Now we solve Eq. (32), looking for $B$ by the perturbation theory in small $\beta$; for that we use the expression Eq.(33) while calculating the factor $B^{3/2}$. We consider the case $\Delta,Pk_{y}\ll E_{g}$. Then for quantity $B_{0}\equiv B|_{\beta=0}$ it is enough to use a simple expression $B_{0}=P^{2}k_{y}^{2}/2E_{g}$. Using the notation $x=P^{2}k_{y}^{2}/2B$, and assuming $B\ll\Delta$ (which actually means that $Pk_{y}\ll\sqrt{E_{g}\Delta}$), we obtain the desired spectrum of the DK1 surface branch
$x$ $=$ $E_{g}-\Delta/2+\Delta(1-\tilde{k}_{y}/|k_{y}|)(\tilde{k}_{y}/|k_{y}|-1-2k_{y}/k_{\Delta}),$
$\epsilon_{DK1,\downarrow}$ $=$ $\frac{E_{g}}{2}-\frac{\Delta}{2}+\frac{P^{2}k_{y}^{2}}{2x}$ (35)
Here $\tilde{k}_{y}=\sqrt{2\beta}E_{g}/P,\,\,k_{\Delta}=\sqrt{6}\Delta/P$. We will consider the case of small $\beta\ll(\Delta/E_{g})^{2}\ll 1$, then two characteristic momenta $\tilde{k}_{y}$ and $k_{\Delta}$ relate as $\tilde{k}_{y}\ll k_{\Delta}$. The first one, $\tilde{k}_{y}$, is the cut-off value at which the spin-down DK1 branch starts splitting off the HH band, see Fig. 2. (At $k_{y}<\tilde{k}_{y}$ the quantity $\kappa_{l}$, Eq. 32, is not real). We remind that for $\beta=0$ case the starting point was zero $k_{y}$ value Khaetskii . At large values of $k_{y}$ this branch is described by the first line of Eq. (33).
The second branch of surface states in this energy interval near the $\Gamma_{8}$ point corresponds to the spin-up state and is obtained by changing the sign of $k_{y}$ in Eq. (35). This branch (at $k_{y}>0$ ) also starts near $\tilde{k}_{y}$ (with slight shift from this point determined by small quantity $\tilde{k}_{y}/k_{\Delta}\ll 1$), has nonmonotonic behavior, and merges again with the heavy-hole band at the point $k_{\Delta}$, Fig. 2. This branch exists within the Kane model already in the case of $\beta=0$, see the second line of Eq. (33). Since the interval of $k_{y}$-values at which it shows up is narrow and is determined by a small strain value, we did not describe this branch in our previous work, Ref. Khaetskii, . (Note that within the Luttinger model, when $E_{g}\to\infty$, this spin-up branch (at $k_{y}>0$) disappears and Fig. 2 reduces to Fig. 6(b) of Ref. Khaetskii, ). Thus, based on the results obtained here (Fig. 2), we can exclude the formation of a twin Dirac cone for the system we consider.
@@
Figure 2: Surface states around the $\Gamma_{8}$ point for a semi-infinite sample within the Kane model with a finite HH mass and positive strain, $\Delta>0$. Blue dots correspond to the cut-off $\tilde{k}_{y}$ points where the DK1 surface branch splits off the HH band, see Eq. (35). Black dots correspond to the $k_{\Delta}$ points. Surface branches are nondegenerate. The spin orientations are shown by the red and blue arrows.
V Gap in the V-P states for the finite-size sample

### v2

@@ Abstract
We consider surface states in semiconductors with inverted-band structures, such as $\alpha$-Sn and HgTe. The main interest is the interplay of the effect of a strain of an arbitrary sign and that of the sample finite size. We consider, in particular, a model system comprised of a gapless semiconductor (e.g. HgTe or $\alpha$-Sn) of finite-width sandwiched between layers of a regular-band semiconductor (e.g. CdTe or InSb). We clarify the origin of various transitions that happen at a given strain with the change of the sample thickness, in particular the transition between the Dirac semimetal and quasi-3D (quantized) topological insulator. Our conclusion opposes those reached recently by the majority of researchers. We show that near the transition point the surface state cannot be treated as a truly topological one since the parameters of the problem are such that an appreciable overlap of the surface states’ wave functions located at opposite boundaries occur. As a result, a spin-conserving, elastic impurity scattering between the states located at opposite boundaries will induce substantial backscattering and destroy the robustness of the surface state . For the k-p Kane model we derive hard-wall boundary conditions in the case when the regular-band materials form high barriers for the carriers of the inner inverted-band semiconductor (for example, CdTe/HgTe/CdTe and CdTe/$\alpha$-Sn/CdTe cases). We show that in this case the boundary conditions have universal and simple form and allow investigation of the realistic case of finite mass of the heavy-hole band, and comparison of the results obtained within the Kane and Luttinger models. In particular, a new type of surface states (wing states) developes with application of strain in the Kane model and is absent in the Luttinger model.
@@ I Introduction
Describing the electronic surface states of gapless semiconductors has long been an interesting problem because of the rich physics at their interfaces. Gapless semiconductors, like HgTe (or $\alpha$-Sn), have inverted band structures where the s-like $\Gamma_{6}$ ($\Gamma_{7-}$) and p-like $\Gamma_{8}$ ($\Gamma_{8+}$) bands switch order in energy due primarily to strong spin-orbit coupling. Groves ; Bloom ; Gelmont Crystalline symmetry enforces a four-fold degeneracy at the $\Gamma_{8}$-point, effectively closing the bandgap. Because of the band inversion, interfaces with vacuum or with direct band-gap semiconductors having ordinary band ordering can produce new types of interface (surface) states, namely Dyakonov-Khaetskii (DK) states Dyak and Volkov-Pankratov (VP) states Volkov ; Cade ; Suris ; Volkov1 . The DK states lie above the $\Gamma_{8}$ heavy-hole band in HgTe and $\alpha$-Sn materials, whereas the VP branch is located below the heavy-hole band within the $\Gamma_{8}-\Gamma_{6}$ gap. These two types of states always coexist since they are just two branches of the same topological Dirac state repelled from each other by the coupling to the heavy-hole band. Moreover, these surface states have nondegenerate spin-texture with a strong spin-momentum locking, making them interesting for spintronics and optoelectronics applications. Ando Although these surface states were predicted in the 1980s, it has been only recently that the topological nature of the states was recognized. Hasan ; Qi ; Bernevig ; Bernevig1 ; Bernevig2 ; Kane ; Dai ; Khaetskii ; Pankratov Experimentally, these interface states can be revealed by angle- and spin-resolved photoemission spectroscopy Xia ; Hsieh and magneto-transport measurements. Brune ; Kozlov ; Checkelsky ; Analytis ; Ren ; Xia1 ; Pan
@@
The energy dispersion of interface states becomes more complex for many practical samples of a thin topological insulator deposited onto a topologically trivial semiconductor or insulator. Under these circumstances, both finite-size effects and epitaxial strain due to a lattice-mismatch between a film of topological material and substrate can significantly modify the electronic band structure, causing transitions between different topological states. Because future devices employing topological materials will likely be made using thin films, the interplay of the finite-size effects and strain within a topological insulator is an important and interesting problem.
In the current literature, there is a broad range of opinions about the origin and exact conditions for topological transitions of thin-film TIs Ohtsubo ; Coster ; Anh ; Vail . Many researchers who primarily use numerical methods do not provide insight into the physical origin of these transitions, because they lack a clear physical model of the bulk-like and surface-like subbands involved in these transitions. In sections V-VII, we discuss our classification of states which is different compared to those made in Refs. Ohtsubo, ; Coster, ; Anh, ; Vail, . This allows one to describe correctly the physics of the transitions in question.
The main goal of this work is to understand the physics of the series of transitions that happen at a given strain value with changing sample width. In particular, we have considered analytically the transition from the Dirac semimetal to the 3D quantized topological insulator that occurs in the presence of a biaxially compressive in-plane strain. In the limit of large width, this strain leads to an overlap of the light- and heavy-hole bands (DSM regime). We show that the DK states are responsible for this transition. Using the exactly solvable Luttinger model, we identify reliably all the bulk-like and surface-like subbands involved in the transition and make concrete predictions about the robustness of the topological states.
@@
We consider here a model system comprised of a gapless semiconductor with an inverted band structure (e.g. HgTe or $\alpha$-Sn) of finite-width $d$ sandwiched between layers of a semiconductor with a regular band order (e.g. CdTe or InSb), see Fig. 1. In section II, we present the Hamiltonian of the Kane model which includes also the contributions from the far bands; thus the heavy-hole (HH) band acquires a finite mass, which allows one to describe the physics applicable for realistic situations. In section III, we consider this model with strain to describe the interface between the inverted-band and direct-band semiconductors focusing on the energies of the DK surface states Dyak ; Khaetskii . In section IV, we derive an exact formula for the location of the Dirac point of the VP states (in the case of a single interface). In section V, we introduce finite-size effects and show how they split the VP states and discuss how it relates to possible interpretations of experimental data (for example, ARPES). In section VI, we describe the interplay between strain and finite-size effects for cases of biaxially compressive and tensile strain. Finally, in section VII, we discuss our results emphasizing their difference and novelty compared to the ones obtained by the other research teams. In Appendix A, we develop the necessary boundary conditions (BCs) for a proper solution under finite thickness. In the case when the direct-band materials form high barriers for the carriers of the inner inverted-band semiconductor, the BCs have an universal and simple form, thereby offering simple analytical treatment and more physical insight than achieved previously.
We stress that our main results address the physics of surface states in the quasi-3D quantum wells (QWs), which differs from the work by Zhou et al. in Ref. Zhou, where the edge states in the 2D QWs were studied. The transition from the 2D normal insulator state to the 2D topological insulator state (2D NI-2D TI) occurs at $\simeq$5.8 nm for the CdTe/$\alpha$-Sn/CdTe structure, whereas the DSM to quasi-3D TI transition addressed in this work occurs at $\simeq$25 nm in thickness. The first case is discussed in Sec. V and the second one is considered in Sec. VI.2. The physics corresponding to these cases is totally different in that they involve the coupling of states in different dimensions. The work by Zhou et al. describes the coupling between the 1D edge channels across the width of a 2D sample near the thickness at which the transition 2D NI-2D TI just happened. It is important to mention that for such small thicknesses, strain does not typically play a role. The physics we consider in Section VI in relation to, for example, the DSM-quasi 3D TI transition is necessarily determined by the strain value, and happens at a significantly greater thickness. Therefore, we consider a quasi-3D sample and study the coupling between the 2D topological surface states.
@@
Figure 1: Left: Configuration of the structure under consideration. A gapless semiconductor (region 1) of thickness $d$ is bounded on parallel sides by a regular semiconductor (region 2). A Cartesian coordinate system is placed such that the $yz$-plane is within one interface where $x=0$. Right: Idealized energy band diagram of the structure. The parameter $\eta$ describes the asymmetry of the offsets in the conduction and valence bands.
@@ II Material Configuration and Hamiltonian
In the end of this section let us make the following important remark. As it follows from Eq. (3), there are only two solutions for $k_{x}^{2}$ of the characteristic equation, i.e. a solution for the heavy holes and a solution for light particles (which are the mixture of electrons and light holes). Therefore, only two independent coefficients exist in the expression for a wave function that is a solution of the Hamiltonian Eq. (2). On the other hand, a spinor $\hat{\Psi}$ has three components. That is why the open boundary condition $\hat{\Psi}=0$ for the case of high barriers is not valid. The correct hard-wall boundary condition is derived in Appendix A.
III Wing states in the Kane model with strain
@@
One of the goals of this work is an attempt to clarify existing ARPES data. In this respect, the surface state spectrum near the $\Gamma_{8}$ energy for the realistic case of the finite heavy-hole mass is interesting because it allows to exclude or confirm the possibility of forming the twinning of Dirac cones,Seixas ; Balatsky which we set out to do here for our particular system.
@@
Two types of surface branches are found within the Kane model. The parabolic branch (called DK1 state) Dyak ; Khaetskii starts from the HH band and ends at the light-hole (LH) band. The Volkov-Pankratov states lie within the $\Gamma_{8}-\Gamma_{6}$ gap Volkov . (These two types of states always coexist since they are just two branches of the original topological Dirac state repelled from each other by the repulsive interaction with the heavy-hole band Khaetskii ). The DK1 branch in the presence of strain was studied in Ref. Khaetskii, for the idealized case of a flat heavy-hole band. Here we investigate this problem at $\Delta>0$ (biaxial, tensile in-plane strain) and $\eta=1/2$ (symmetric band offsets). The parameter $\eta$ ($10<\eta<$ 1) describes the asymmetry of the offsets in the conduction and valence bands, see Fig. 1. We assume a small ratio $\beta$ of the light and heavy hole masses given by $\beta=3\gamma_{h}E_{g}/2P^{2}\ll 1$, $\gamma_{h}=(\gamma-2\tilde{\gamma})$ We will use the perturbation approach with respect to $\beta\ll 1$.
@@
Using the boundary conditions in Eq. (39), one obtains the folowing equation for the surface states spectra
@@
The quantity $B$ in Eqs. (10,13) is defined as $B=\epsilon+\Delta/2-E_{g}/2$. The equation for the time-reversed branch of the states is obtained by replacing $k_{y}$ by $-k_{y}$ in Eq. (13).
@@
where $\epsilon_{HH}=E_{g}/2-\Delta/2$ is the energy of the bottom of the HH band, and $\epsilon_{DP}=\Delta/4$ is the location of the Dirac point of the Volkov-Pankratov states (at $k_{y}=0$). Note that in contrast to the $\Delta=0$ case, for a finite positive strain there are two DK1 branches corresponding to the opposite spins for a given value of $k_{y}$. The one corresponding to the spin-up state (for $k_{y}>0$) exists only up to the critical value of $k_{y}=k_{\Delta}=\sqrt{6}\Delta/P$ (assuming that $\Delta\ll E_{g}$), where it merges with the heavy-hole band, see Fig. 2 and the discussion below. This branch we call the wing state.
For completeness we also present here (at $\beta=0$ ) the exact spectra of two VP surface branches that lie within the $\Gamma_{8}-\Gamma_{6}$ energy gap (and which we derived using the formulas from Ref. Khaetskii, ):
@@
Note that Eq. (15) is the generalization for the case of $\Delta\neq 0$ of the equations of Ref. Volkov1, . Now we solve Eq. (13), looking for $B$ by the perturbation theory in small $\beta$; for that we use the expression Eq.(14) while calculating the factor $B^{3/2}$. We consider the case $\Delta,Pk_{y}\ll E_{g}$. Then for quantity $B_{0}\equiv B|_{\beta=0}$ it is enough to use a simple expression $B_{0}=P^{2}k_{y}^{2}/2E_{g}$. Using the notation $M=P^{2}k_{y}^{2}/2B$, and assuming $B\ll\Delta$ (which actually means that $Pk_{y}\ll\sqrt{E_{g}\Delta}$), we obtain the desired spectrum of the DK1 surface branch
$M$ $=$ $E_{g}-\Delta/2+\Delta(1-\tilde{k}_{y}/|k_{y}|)(\tilde{k}_{y}/|k_{y}|-1-2k_{y}/k_{\Delta}),$
$\epsilon_{DK1,\downarrow}$ $=$ $\frac{E_{g}}{2}-\frac{\Delta}{2}+\frac{P^{2}k_{y}^{2}}{2M}$ (16)
Here $\tilde{k}_{y}=\sqrt{2\beta}E_{g}/P,\,\,k_{\Delta}=\sqrt{6}\Delta/P$. We will consider the case of small $\beta\ll(\Delta/E_{g})^{2}\ll 1$, then two characteristic momenta $\tilde{k}_{y}$ and $k_{\Delta}$ relate as $\tilde{k}_{y}\ll k_{\Delta}$. The first one, $\tilde{k}_{y}$, is the cut-off value at which the spin-down DK1 branch starts splitting off the HH band, see Fig. 2. (At $k_{y}<\tilde{k}_{y}$ the quantity $\kappa_{l}$, Eq. 13, is not real). Recall that for $\beta=0$ case the starting point was zero $k_{y}$ value Khaetskii . At large values of $k_{y}$ this branch is described by the first line of Eq. (14).
Figure 2: Surface states around the $\Gamma_{8}$ point for a semi-infinite sample within the Kane model with a finite HH mass and positive strain, $\Delta>0$. Blue dots correspond to the cut-off $\tilde{k}_{y}$ points where the DK1 surface branch splits off the HH band, see Eq. (16). Black dots correspond to the $k_{\Delta}$ points, where the wing states terminate. Surface branches are nondegenerate. The spin orientations are shown by the red and blue arrows.
The second (wing) branch of surface states in this energy interval near the $\Gamma_{8}$ point corresponds to the spin-up state and is obtained by changing the sign of $k_{y}$ in Eq. (16). This branch (at $k_{y}>0$ ) also starts near $\tilde{k}_{y}$ (with slight shift from this point determined by small quantity $\tilde{k}_{y}/k_{\Delta}\ll 1$), has nonmonotonic behavior, and merges again with the heavy-hole band at the point $k_{\Delta}$, Fig. 2. This wing branch exists within the Kane model already in the case of $\beta=0$, see the second line of Eq. (14). Since the interval of $k_{y}$-values at which it shows up is narrow and is determined by a small strain value, we did not describe this branch in our previous work, Ref. Khaetskii, . (Note that within the Luttinger model, when $E_{g}\to\infty$, the wing branches disappear and Fig. 2 reduces to Fig. 6(b) of Ref. Khaetskii, ). Thus, based on the results obtained here (Fig. 2), we can exclude the formation of a twin Dirac cone for the system we consider.
@@
IV Location of the Dirac point of the Volkov-Pankratov states
Existing ARPES data reveal two experimental facts that strongly disagree with the theoretical predictions of many papers. Some ARPES experiments Engel ; Polaczynski with $\alpha$- Sn material show that the DP lies in close proximity to the $\Gamma_{8}$ point. From Eq. (17) it follows that for the actual parameters of the structures which have the energy diagram of the type presented in Fig. 1 this fact cannot be explained. The other puzzling experimental fact is that the dispersion of the surface states (presumably of the VP types) appears linear in k-vector and connect s-type band with the light-hole band crossing the fundamental gap without experiencing any repulsive interaction with the bulk HH band. The last fact contradicts the theoretical predictions, see, for example, Ref. Khaetskii, .
For arbitrary values of the gap energies, $E_{g}>0,E_{gb}>0$, in the inverted and direct (barrier) materials, and arbitrary offsets in the conduction and valence bands, we obtain the following formula for the location of the Dirac point of the Volkov-Pankratov states Volkov
A similar result was obtained earlier in Ref. Cade, . The zero of the energy coincides with the middle of the $\Gamma_{8}-\Gamma_{6}$ gap of the inverted material. While deriving Eq. (17) we neglected the $\gamma,\tilde{\gamma}$ terms in the Hamiltonian and used the continuity of the electron and light-hole components of spinors found at $k_{y}=0$ in the inverted-band and direct-gap materials.
The location of the DP close to the $\Gamma_{8}$ point can be obtained from Eq. (17) in the case when $E_{gb}\gg E_{g}$ and $\eta\to 1$. $\eta\to 1$ means that the offset in the conduction band is much larger than in the valence band. Then at $\Delta=0$ one obtains $\epsilon_{DP}=E_{g}/2$, i.e. at the energy of the $\Gamma_{8}$ point of the inverted material.
If, however, one considers the case of $\alpha$-Sn/InSb interface Engel , then using the values of offsets from Ref. Cardona, and assuming a zero strain, one obtains the location of the DP to be rather closer to the middle of the $\alpha$-Sn gap, which differs from the interpretations Engel of the ARPES measurements. It should be mentioned that the situation does not become better even if an interaction with the far bands is taken into account. In Appendix B the correction to Eq. (17) due to the terms proportional to $\gamma$ and $\tilde{\gamma}$ in the Hamiltonian Eq. (2) is derived. We have shown that such correction is negative for the cases we consider in this work (HgTe/CdTe and $\alpha$-Sn/InSb), i.e. a shift of the DP is downward in energy with respect to the location given by Eq. (17) and cannot explain the experimental data.
We suggest both experimental facts mentioned in the beginning of this section cannot be explained within the idealized theoretical model presented by Fig. 1. Recall that for the typical range of ARPES excitation energies, the escape depth of electrons is about 1 nm near the interface. Here we refer to the so-called ’universal curve” for all electron spectroscopies including ARPES. Seah_Dench Consequently, information about the surface states is collected only within $\sim$ 1 nm near the vacuum surface. If a bending of the energy bands near the interface in the inverted-band material (HgTe or $\alpha$-Sn) occurs in such a way that it creates a potential barrier for the heavy holes, then they just cannot reach the interface layer which is probed by ARPES. Perhaps only the heavy holes that travel exactly perpendicular to the surface can tunnel into the interface layer, but they are decoupled from the light particles anyway Khaetskii . Under such conditions the ARPES pattern will look like in Fig. 1 of Ref. Khaetskii, . Moreover, the picture of the band bending described above means that the energy of the $\Gamma_{8}$ point in the inverted-band material is shifted downward, and the DP of the VP states may be close to the $\Gamma_{8}$ point.
V Gap in the Volkov-Pankratov states